40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3256     Entry time: Wed Jul 21 12:03:14 2010
Author: Katharine, Sharmila 
Type: Update 
Category: WIKI-40M Update 
Subject: weekly update 

This past week, we levitated our small cylindrical magnet (with the flag made from heat shrink).  Though the levitated magnet didn't appear very jittery to the eye, we looked at the PD current on the scope and could see oscillations that corresponded to the flag hitting the sides of the OSEM.  The oscillations were more pronounced as we gently hit/vibrated the lab bench, and by pounding on the bench Rana knocked the levitated magnet completely out of the setup.  So, we're currently working on building a new, stabler mount.  The biggest challenge here is fixing the OSEM in place, but we're experimenting with different optics pieces to see which is the most stable for our purposes.   Jenne taught us how to make through holes using the drill press so we can add slats of aluminum to adjust the height of the OSEM mount.  We also plan to fix some heavy plates to the bottom of our system to decrease its vibration frequency.

We also calculated the transfer function of our circuit, which seems to match the measured frequency response to within a small factor.  We're playing with Rana's Simulink models and are currently modeling our own system to determine what gains we would expect to use and get a better understanding of our circuit.


Once our system is successfully mounted stably, we plan to experimentally observe the effects of changing the gain and integrator by looking at time series measurements of the PD current, as well as using the spectrum analyzer to compare the frequency response of our system at different gain settings.

ELOG V3.1.3-